# PromptEnhancer: A Simple Approach to Enhance Text-to-Image Models via Chain-of-Thought Prompt Rewriting

### v1

@@ 2 Methodology
Figure 2: An overview of the training framework for PromptEnhancer. Our framework trains a universal Rewriter to enhance pretrained Text-to-Image (T2I) model without altering its weights. This is achieved through a two-stage process guided by a specialized reward model. Stage 1: SFT for Rewriter Initialization (Sec 2.2). The CoT Rewriter is first initialized via SFT on (user prompt, reprompt) pairs. This stage teaches the model to generate structured, chain-of-thought style responses using a standard next-token prediction loss, establishing a strong foundation for refinement. Stage 2: Policy Alignment with GRPO (Sec 2.3). The initialized rewriter is further refined using GRPO. The rewriter generates multiple reprompt candidates, which are used by a frozen T2I model to create images. The pre-trained AlignEvaluator then assesses each (image, reprompt) pair and provides a scalar reward. This reward signal optimizes the rewriter's policy, steering it toward generating prompts that maximize image-text alignment. The AlignEvaluator (Sec 2.4). Central to our framework is the AlignEvaluator, a pre-trained reward model. It is trained on a large-scale dataset annotated against a taxonomy of 24 fine-grained key points (T2I-KeyPoints, Tab. 1). This enables it to provide a robust and nuanced reward signal, which is crucial for the policy alignment stage.
@@ 2.3 Stage 2: Policy Alignment with GRPO
Policy Alignment via GRPO. After SFT, the rewriter can generate plausible long prompts, but they are not explicitly optimized for image-text alignment. The second stage therefore employs Group Relative Policy Optimization (GRPO (Shao et al., 2024)) to align the rewriter's output with the fine-grained preferences captured by our AlignEvaluator.
The GRPO Training Loop. This training process is iterative. For a given prompt, the CoT Rewriter first generates $N$ candidate rewritten prompts $\{p^{\prime}_{1},p^{\prime}_{2},...,p^{\prime}_{N}\}$. Each of these prompts is then fed into the frozen T2I model to generate a corresponding image $I_{i}$. Subsequently, our AlignEvaluator calculates a scalar reward $r_{i}$ for each pair $(p^{\prime}_{i},I_{i})$. Finally, these rewards $\{r_{1},r_{2},...,r_{N}\}$ are used to update the rewriter's policy, creating a preference ranking that encourages the generation of prompts leading to higher rewards.
@@ 3.1 SFT Data for Rewriter Initialization
The dataset is balanced between English and Chinese. Thematically, it primarily covers general scenarios (approx. 90%), with the remainder dedicated to specialized domains such as intellectual property (IP) characters (4%), in-image text rendering (3%), and specific camera viewpoints (3%).
1. User Prompt Simulation. To emulate concise user queries at scale, we begin with a diverse pool of 3.26 million images (1.53M Chinese-centric and 1.73M Western-centric). We employ an image captioning model to generate short, naturalistic descriptions for these images. This step results in a collection of 2.26 million proxy user prompts that serve as the foundation for our data generation.
@@
3. Human-in-the-Loop Selection. To ground our dataset in perceptual quality, we introduce a critical human evaluation step. For each set of generated reprompt candidates, we use the Hunyuan Text-to-Image model to synthesize the corresponding images. Professional annotators then assess these images, selecting the single reprompt that best aligns with the user's intent and yields the highest-quality visual output. This human-in-the-loop process produced a curated set of 611,921 triplets.
4. Automated Filtering. The final stage involves a rigorous automated filtering process to further enhance data quality. We again use Gemini-2.5-Pro to programmatically detect and discard samples with issues such as semantic deviation, information loss, or linguistic incoherence. This step refines the dataset from 611,921 instances to a final corpus of 485,119 triplets, ensuring a high degree of fidelity for model training.
@@ 3.2 Prompt Set for Policy Alignment
Figure 7: End-to-End Workflow and Impact of PromptEnhancer. This figure illustrates the complete process and impact of the PromptEnhancer framework across four different use cases. Each case is presented in a four-column format: (1) the initial user prompt, (2) the baseline image generated from that prompt, (3) the detailed, rewritten prompt from the CoT rewriter, and (4) the final, superior image generated from the rewritten prompt. This clearly shows the framework's ability to handle complex scenes involving in-image text (e.g., ``Drive the Future") and specific layouts (e.g., an Instagram Reels cover), which are common challenges for text-to-image models.
Figure 8: Comparative Analysis Against State-of-the-Art Models. This figure provides a head-to-head comparison of PromptEnhancer against several other leading text-to-image models, including GPT-4o (Hurst et al., 2024), Seedream 3.0 (Gao et al., 2025), Kolor 2.1 (Team, 2024), and Qwen-Image (Wu et al., 2025a), as well as the baseline Hunyuan Image 2.1 model. Across a variety of challenging prompts that require accurate text rendering and compositional integrity, the images generated using PromptEnhancer consistently demonstrate superior prompt adherence. While other models produce aesthetically pleasing results, they often struggle with fine-grained details, leading to common errors such as garbled text or misinterpreted scene layouts.
@@ 4.3 Quantitative Analysis
As illustrated in Figure 6, our framework provides a significant and broad-spectrum enhancement to the T2I model's prompt-following capabilities. On average, PromptEnhancer boosts the accuracy across all 24 evaluation dimensions by a substantial 5.1 percentage points (pp), elevating the mean score from a baseline of 65.9% to 71.0%. This overall improvement is underpinned by widespread gains, with performance increasing in 21 out of 24 categories.
Significant Gains in Complex Reasoning. The most pronounced improvements are concentrated in categories that demand sophisticated semantic and compositional understanding. Notably, we observe dramatic boosts in Similarity Relation (+17.3 pp), Counterfactual reasoning (+17.2 pp), and Counting (+15.0 pp). These results underscore our method's ability to help the T2I model interpret abstract relationships and precise numerical constraints. Furthermore, significant enhancements are seen in fine-grained linguistic and visual challenges, including Pronoun Resolution (+13.9 pp), rendering nuanced facial Expression (+12.9 pp), and maintaining Cross-Entity Binding (+11.3 pp). In total, 15 distinct categories exhibit an accuracy increase of more than 5 pp, demonstrating the comprehensive impact of our prompt rewriting strategy.
Areas of Neutrality and Minor Regression. While the impact is overwhelmingly positive, the gains are not uniform. Performance in Contact Interaction remains unchanged (+0.0 pp), and Artistic Style shows only a marginal improvement (+0.9 pp), suggesting the baseline model is already relatively competent in these areas. We identify minor performance regressions in three categories: Interaction w/o Contact (-4.9 pp), Size (-0.7 pp), and Text Layout (-0.7 pp). These isolated cases represent valuable areas for future work, indicating that for certain simple concepts, the rewriting process might occasionally introduce ambiguity or over-specification. Nevertheless, these minor trade-offs are far outweighed by the extensive and significant improvements across the vast majority of challenging categories.
4.4 Qualitative Analysis
A qualitative review of the generated images confirms that the PromptEnhancer framework provides a significant advantage in enhancing the detail, accuracy, and artistic quality of text-to-image outputs. The core strength of the framework lies in its ability to automatically convert a user's concise prompt into a highly detailed and structured description through a Chain-of-Thought process (Figure 5, Figure 7). This enrichment process adds specific entity details (e.g., clarifying a ``cute cat" as ``Tom cat from Tom & Jerry"), precise artistic styles (e.g., ``oil painting texture"), and exact compositional instructions, which vastly improves the depth and quality of the final image.
Furthermore, the framework demonstrates exceptional performance in handling complex scenes, particularly in rendering in-image text, which is a notorious challenge in the field. As shown in the side-by-side comparisons, PromptEnhancer successfully guides the model to generate clear and accurate text for billboards and social media posts, such as ``Drive the Future'' and ``PawPerfect'' (Figure 7 and Figure 8). In a direct comparison against other models, PromptEnhancer consistently exhibits superior robustness in faithfully adhering to these complex instructions. While other models often produce images with garbled text or incorrect compositions, this framework yields results that are more precisely aligned with the user's specific and nuanced requests (Figure 8).
@@ 5.1 Text-to-Image Generation
The field of Text-to-Image (T2I) synthesis has witnessed a paradigm shift with the advent of diffusion models (ho2020denoising; song2020score), which have largely superseded earlier GAN-based approaches (Reed et al., 2016). Seminal models like DALL-E (Ramesh et al., 2021), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022) established a powerful foundation for generating high-fidelity and diverse images. This frontier continues to be pushed by state-of-the-art systems such as DALL-E 3 (Betker et al., 2023), and more recent architectures like FLUX.1 (Labs et al., 2025), which have dramatically improved photorealism and text rendering.
However, despite this progress, a fundamental challenge persists: the prompt alignment gap. Models often struggle with compositional complexity, failing to correctly bind attributes to objects, comprehend spatial or logical relationships (e.g., negation), and maintain fidelity to intricate details (Huang et al., 2023). This gap places a significant "alignment tax" on users, forcing them to engage in tedious and often unintuitive prompt engineering. Our work directly targets this gap, not by inventing a new generator, but by creating a universal interface that teaches any T2I model to better understand human intent.

### v2

@@ 2 Methodology
Figure 2: An overview of the training framework for PromptEnhancer. Our framework trains a universal Rewriter to enhance pretrained Text-to-Image (T2I) model without altering its weights. This is achieved through a two-stage process guided by a specialized reward model. Stage 1: SFT for Rewriter Initialization (Sec 2.2). The CoT Rewriter is first initialized via SFT on (user prompt, reprompt) pairs. This stage teaches the model to generate structured, chain-of-thought style responses using a standard next-token prediction loss, establishing a strong foundation for refinement. Stage 2: Policy Alignment with GRPO (Sec 2.3). The initialized rewriter is further refined using GRPO. The rewriter generates multiple reprompt candidates, which are used by a frozen T2I model to create images. The pre-trained AlignEvaluator then assesses each (image, user prompt) pair and provides a scalar reward. This reward signal optimizes the rewriter's policy, steering it toward generating prompts that maximize the alignment between the image and the user's intent. The AlignEvaluator (Sec 2.4). Central to our framework is the AlignEvaluator, a pre-trained reward model. It is trained on a large-scale dataset annotated against a taxonomy of 24 fine-grained key points (T2I-KeyPoints, Tab. 1). This enables it to provide a robust and nuanced reward signal, which is crucial for the policy alignment stage.
@@ 2.3 Stage 2: Policy Alignment with GRPO
Policy Alignment via GRPO. After SFT, the rewriter can generate plausible long prompts, but they are not explicitly optimized for image-text alignment. The second stage therefore employs Group Relative Policy Optimization (GRPO (Shao et al., 2024)) to align the user prompt with the fine-grained preferences captured by our AlignEvaluator.
The GRPO Training Loop. This training process is iterative. For a given user prompts $\{p_{1},p_{2},...,p_{N}\}$, the CoT Rewriter first generates $N$ candidate rewritten prompts $\{p^{\prime}_{1},p^{\prime}_{2},...,p^{\prime}_{N}\}$. Each of these prompts is then fed into the frozen T2I model to generate a corresponding image $I_{i}$. Subsequently, our AlignEvaluator calculates a scalar reward $r_{i}$ for each pair $(p_{i},I_{i})$. Finally, these rewards $\{r_{1},r_{2},...,r_{N}\}$ are used to update the rewriter's policy, creating a preference ranking that encourages the generation of prompts leading to higher rewards.
@@ 3.1 SFT Data for Rewriter Initialization
Based on fig. 4, the dataset exhibits a thematic distribution across several creative domains. Design constitutes the largest portion at 27%, followed by Art at 23%, and Film & Story at 22%. Illustration accounts for 18%, and Creative tasks make up 10% of the dataset.
1. User Prompt Simulation. To emulate concise user queries at scale, we begin with a diverse pool of 3.26 million images (1.53M Chinese-centric and 1.73M English-centric). We employ an image captioning model to generate short, naturalistic descriptions for these images. This step results in a collection of 2.26 million proxy user prompts that serve as the foundation for our data generation.
@@
3. Automated Filtering. The initial stage involves a rigorous automated filtering process aimed at enhancing data quality. We utilize Gemini-2.5-Pro to programmatically detect and discard samples with issues such as semantic deviation, information loss, or linguistic incoherence. This process reduces the dataset from 1 million instances to 611,921 triplets.
4. Human-in-the-Loop Selection. To further ground our dataset in perceptual quality, we introduce a critical human evaluation step. For each set of generated reprompt candidates, we use the Hunyuan Text-to-Image model to synthesize corresponding images. Professional annotators then evaluate these images, selecting the reprompt that best aligns with the user's intent and produces the highest-quality visual output. This human-in-the-loop process results in a final curated set of 485,119 triplets.
@@ 3.2 Prompt Set for Policy Alignment
Figure 4: Distribution of Categories in the Dataset. The chart on the left shows the primary categories, while the chart on the right provides a detailed breakdown into 20 sub-categories.
@@ 4.3 Quantitative Analysis
As illustrated in Figure 7, our framework provides a significant and broad-spectrum enhancement to the T2I model's prompt-following capabilities. On average, PromptEnhancer boosts the accuracy across all 24 evaluation dimensions by a substantial +5.1%. This overall improvement is underpinned by widespread gains, with performance increasing in 20 out of 24 categories.
Significant Gains in Complex Reasoning. The most pronounced improvements are concentrated in categories that demand sophisticated semantic and compositional understanding. Notably, we observe dramatic boosts in Similarity Relation (+17.3%), Counterfactual reasoning (+17.2%), and Counting (+15.0%). These results underscore our method's ability to help the T2I model interpret abstract relationships and precise numerical constraints. Furthermore, significant enhancements are seen in fine-grained linguistic and visual challenges, including Pronoun Resolution (+13.9%), rendering nuanced facial Expression (+12.0%), and maintaining Cross-Entity Binding (+11.3%). In total, 15 distinct categories exhibit an accuracy improvement of more than 5.0%, demonstrating the comprehensive impact of our prompt rewriting strategy.
Areas of Neutrality and Minor Regression. While the impact is overwhelmingly positive, the gains are not uniform. Performance in Contact Interaction and Size remains unchanged (+0.0%), and Artistic Style shows only a marginal improvement (+0.9%), suggesting the baseline model is already relatively competent in these areas. We identify minor performance regressions in two categories: Text Layout (-0.7%) and Interaction w/o Contact (-0.9%). These isolated cases represent valuable areas for future work, indicating that for certain simple concepts, the rewriting process might occasionally introduce ambiguity or over-specification. Nevertheless, these minor trade-offs are far outweighed by the extensive and significant improvements across the vast majority of challenging categories.
@@ 5.1 Text-to-Image Generation
The field of Text-to-Image (T2I) synthesis has witnessed a paradigm shift with the advent of diffusion models (Ho et al., 2020; Song et al., 2021), which have largely superseded earlier GAN-based approaches (Reed et al., 2016). Seminal models like DALL-E (Ramesh et al., 2021), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022; Podell et al., 2024; Esser et al., 2024) established a powerful foundation for generating high-fidelity and diverse images. This frontier continues to be pushed by state-of-the-art systems such as DALL-E 3 (Betker et al., 2023), and more recent architectures like FLUX.1 (Labs et al., 2025), which have dramatically improved photorealism and text rendering.
However, despite this progress, a fundamental challenge persists: the prompt alignment gap. Models often struggle with compositional complexity, failing to correctly bind attributes to objects, comprehend spatial or logical relationships (e.g., negation), and maintain fidelity to intricate details (Huang et al., 2023). This gap places a significant ``alignment tax'' on users, forcing them to engage in tedious and often unintuitive prompt engineering. Our work directly targets this gap, not by inventing a new generator, but by creating a universal interface that teaches any T2I model to better understand human intent.